OpenStudy (anonymous):

factor completely 5x^2-20y^2

OpenStudy (dan815):

difference of squares

OpenStudy (anonymous):

What?

OpenStudy (dan815):

\[(\sqrt{5}x+\sqrt{20}y)(\sqrt{5}x-\sqrt{20}y), Exapand\quad me\]

OpenStudy (anonymous):

I don't know what that is...

OpenStudy (dan815):

\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

So.. 5x^2-20y^2=(5x^2+20y^2)(5x^2-20y^2)?

OpenStudy (dan815):

Do you know how to expand brackets, or foil method?

OpenStudy (dan815):

Have you ever factored before?

OpenStudy (anonymous):

I've never heard of expanding brackets but I can FOIL. No, I have never factored before.